# Graph Neural Networks for Next-Generation-IoT: Recent Advances and Open Challenges

### v1

@@ II-2 State-of-the-art:
Table V provides a summary of existing applications of GNNs on various IoT network scenarios with three types of tasks in the graph, highlighting their effectiveness in solving different types of wireless network challenges.
TABLE V: Existing application of GNNs on various wireless networks categorized by types of tasks
@@
  • Node level: Node-level GNN applications are widely used in wireless networks. One of the most well-known applications of node-level tasks is power allocation, a critical aspect of network optimization [34, 35, 39]. For instance, in [20], the uplink power allocation of a cell-free massive MIMO IoT System can be formulated as a node-level task of an undirected graph, where users and access points (APs) are represented as nodes. This approach allows the model to generalize effectively across different network configurations, ensuring scalability. However, in downlink power allocation, the problem becomes more complex because each AP must allocate power vectors that are dependent on the number of served users, making the task more sensitive to changes in network size.
    To overcome this challenge, Shen et al. [125] proposes an approach, where each communication link is represented as a node. The illustration of communications links as nodes is shown in Fig. 7. This representation shifts the focus from the APs to the links, allowing the system to maintain scalability even as the network size varies. This method ensures that the power allocation process remains efficient and adaptable in dynamic network environments, where the number of connected devices may fluctuate. Furthermore, by representing each communication link as a node, the node-level tasks lend themselves to solving link scheduling problems, such as AP-IoT device association or frequency assignment problems. In this way, decisions regarding transceiver pair scheduling can be made independently based on the node’s features. This method enables efficient link scheduling, as demonstrated in studies like [36, 37].
@@
    Moreover, node-level tasks are particularly effective in predicting user behavior by analyzing the node attributes that capture specific user activity and communication patterns. Since each user’s behavior can be anticipated independently of others, focusing on node-level representations facilitates accurate forecasting of traffic patterns based on local information. This approach is pursued in [40], where a cellular traffic network is modeled as an undirected graph, with mobile traffic data embedded as node attributes. By focusing on these node-level details, the GNN can accurately predict traffic fluctuations and user behavior, leading to more responsive and efficient network management.
@@
    Node-level tasks are also beneficial in classification problems within wireless networks. In [41], a GNN was applied to a binary classification problem to detect anomalies and insider threats. In this context, nodes represent activity log positions, while edges represent the contextual relationships between activities. This node-centric approach allowed the GNN to identify abnormal patterns effectively. Similarly, Huoh et al. [42] addressed encrypted network traffic classification by mapping each packet to a node and using packet relationships and meta-features as graph inputs. The authors represented the classification problem by both a node-level task as well as an edge-level task and revealed that the classification accuracy was higher when using node-level tasks than edge-level tasks. This indicates the effectiveness of node-level representations in certain wireless network applications, where capturing and leveraging node-specific information is crucial for achieving superior performance.
@@
  • Edge level: Edge-level GNNs have proven effective in addressing various wireless network challenges, especially in optimizing user association, which is crucial for enhancing system performance in wireless networks [126]. Liu et al. [43] utilized the edge-GNN concept to predict the user-BS association. Similarly, in [44], the edge-level concept was used to represent the task offloading problem between wireless devices and mobile edge computing. In terms of resource allocation, the edge-level task formulation enhances interference management by directly modeling interactions between interfering links. In dense wireless networks, where devices compete for resources, this approach allows GNNs to predict and mitigate interference more accurately. By focusing on the edges, GNNs can optimize scheduling and power control, hence reducing interference and improving network performance. This is achieved by dynamically adjusting the transmission power in overlapping cellular networks to boost the overall throughput and user experience [45, 46]. Peng et al. [45] revealed that edge-GNN could perform just as well as the node-GNNs in power allocation, with the added benefit of reduced training time.
    Another advantage of edge-level tasks in GNNs is their superior scalability in large and dynamic wireless networks. By focusing on the edges, GNNs can efficiently manage changes in network topology, including the addition or removal of links, without being constrained by the number of nodes. By contrast, node-level tasks may struggle to attain scalability, particularly in scenarios like downlink power allocation. Here, the output dimension at an AP node is typically fixed based on the number of users, limiting flexibility as the network grows. By associating the power allocation variable with the edge, as demonstrated in [46, 47], edge-level GNNs can more effectively adapt to varying network sizes, ensuring scalability.
@@
    Model architecture: We consider the uplink power allocation for a cell-free massive MIMO system. The goal is to maximize the minimum user rate across the system, ensuring efficient energy usage while maintaining high data rates. The max-min fairness criterion harnessed for power allocation is formulated as in [127]. This is particularly challenging, because the users experience varying levels of interference and channel conditions. Efficient power control is crucial for enhancing the system performance attained and for ensuring fairness among users. The global solution may indeed be obtained, but it is computationally expensive, especially for large-scale networks experiencing dynamic user and channel conditions. This makes real-time optimization challenging. Using GNN provides an efficient design alternative.
    As suggested by Peng et al. [45], the cell-free massive MIMO system can be represented as a heterogeneous graph. Therein, the associated max-min fairness problem can be formulated either as an edge-level task or a node-level task and then solved using the edge and node convolution-based methods, respectively. To evaluate the effectiveness of different GNN architectures for this task, we harnessed both models from [45] and proposed new ones. Below is a summary of the models considered:
    • Node-GNN: Focuses on updating node features, representing users and access points, to allocate power based on node-level characteristics.
    • Hybrid Quantum GNN (Our proposal): The HQGNN uses the GNN to preprocess and generate node embeddings, which are then handled by the deep quantum neural network (DQN) using quantum circuits for power allocation.
    • Edge-GNN: Updates edge features, representing the connections between users and access points to capture the relationships between them.
    • Graph-GNN: Utilizes global graph pooling to aggregate features across the entire graph, paving the way for a holistic approach to power allocation.
@@
    Simulations and results: In this simulation, we consider a cell-free massive MIMO system supporting $6$ users with the aid of $30$ access points. Only large-scale fading coefficients are used as inputs of the GNNs. We model the system as a bipartite graph associated with two types of nodes, including cellular users and access points. Fig. 8 presents the performance of various GNN architectures, illustrated using the cumulative distribution function (CDF) derived from $10,000$ realizations of a cell-free massive MIMO system.
    The results show that edge-GNN and node-GNN constitute a pair of models capable of approximating the optimal solution. Briefly, these two models achieve a minimum rate above $1$bps/Hz in about $65$% of the instances. The hybrid quantum GNN, which integrates quantum circuits into the Node-GNN architecture, also performs competitively. Quantitatively, this model tends to achieve $1$bps/Hz in about $64$% of the instances. On average, the Edge-GNN, Node-GNN, and HQGNN models provide data throughputs of $1.0783$, $1.0733$, and $1.0640$ bps/Hz, respectively. By contrast, the optimal solution achieves an average throughput of $1.129$ bps/Hz, slightly outperforming all GNN models, but at the cost of significantly higher computational complexity, particularly as the network size escalates. The graph-GNN performs worst of all GNN models, with less than $40$% of cases reaching a minimum rate of $1.0$ bps/Hz. This poor performance is attributed to its global aggregation of information, which fails to capture the crucial local node and edge interactions necessary for optimizing power allocation in the network. Overall, the results underscore the superiority of node-level GNN models, particularly of those utilizing attention mechanisms for solving the power control problem of cell-free massive MIMO systems. Furthermore, the hybrid quantum GNN offers an intriguing avenue for integrating quantum computing into GNN architectures.
@@ II-3 Challenges and future directions
Again, a significant challenge in modeling GNNs for wireless networks is ensuring scalability, particularly when dealing with networks that have inherent constraints, such as limited node resources or dynamic topology changes. As the graph size grows, the computational cost and memory usage increase, making it challenging to maintain the target performance in real-time. To address this challenge, a promising future direction is to partition the large graph into multiple subgraphs, allowing GNNs to learn locally within each subgraph. The information learned from each subgraph may then be aggregated, ensuring that the model is able to process graphs of varying sizes without losing generalization capabilities. This approach can mitigate scalability issues and maintain performance across different network scales.

### v2

@@ II-2 State-of-the-art:
Table VI provides a summary of existing applications of GNNs on various IoT network scenarios with three types of tasks in the graph, highlighting their effectiveness in solving different types of wireless network challenges.
TABLE VI: Existing application of GNNs on various wireless networks categorized by types of tasks
@@
  • Node level: Node-level GNN applications are widely used in wireless networks, especially in power allocation [34, 35, 39]. For instance, in [20], the uplink power allocation of a cell-free massive MIMO IoT System can be formulated as a node-level task of an undirected graph, where users and access points (APs) are represented as nodes. This approach allows the model to generalize effectively across different network configurations, ensuring scalability. However, in downlink power allocation, the problem becomes more complex because each AP must allocate power vectors that are dependent on the number of served users, making the task more sensitive to changes in network size. To overcome this challenge, Shen et al. [152] proposes to represent communication links as nodes, as illustrated in Fig. 7. This representation shifts the focus to the links, allowing the system to maintain scalability even as the network size varies. This method ensures that the power allocation process remains efficient and adaptable in dynamic network environments, where the number of connected devices may fluctuate. Furthermore, by representing each communication link as a node, the node-level tasks lend themselves to solving link scheduling problems, such as AP-IoT device association or frequency assignment problems. In this way, decisions regarding transceiver pair scheduling can be made independently based on the node’s features. This method enables efficient link scheduling, as demonstrated in studies like [36, 37].
@@
    Node-level tasks also effectively predict user behavior by analyzing node attributes that capture activity and communication patterns. In [40], a cellular traffic network is modeled as an undirected graph with mobile traffic data as node attributes. This approach enables GNNs to forecast traffic fluctuations accurately, enhancing network responsiveness and efficiency. Moreover, node-level tasks are beneficial in classification problems within wireless networks. In [41], a GNN was applied to a binary classification problem to detect anomalies and insider threats. In this context, nodes represent activity log positions, while edges represent the contextual relationships between activities. This node-centric approach allowed the GNN to identify abnormal patterns effectively. Another example, Huoh et al. [42] addressed encrypted network traffic classification by mapping each packet to a node and using packet relationships and meta-features as graph inputs. The authors represented the classification problem by both a node-level task as well as an edge-level task and revealed that the classification accuracy was higher when using node-level tasks than edge-level tasks. This indicates the effectiveness of node-level representations in certain wireless network applications, where capturing and leveraging node-specific information is crucial for achieving superior performance.
@@
  • Edge level: Edge-level GNNs have proven effective in addressing various wireless network challenges, especially in optimizing user association, which is crucial for enhancing system performance in wireless networks [153]. Liu et al. [43] utilized the edge-GNN concept to predict the user-BS association. Similarly, in [44], the edge-level concept was used to represent the task offloading problem between wireless devices and mobile edge computing. In terms of resource allocation, the edge-level task formulation enhances interference management by directly modeling interactions between interfering links. In dense wireless networks, where devices compete for resources, this approach allows GNNs to predict and mitigate interference more accurately. By focusing on the edges, GNNs can optimize scheduling and power control, hence reducing interference and improving network performance [45, 46]. Peng et al. [45] revealed that edge-GNN could perform just as well as the node-GNNs in power allocation, with the added benefit of reduced training time. Another advantage of edge-level tasks in GNNs is their superior scalability in large and dynamic wireless networks. By focusing on the edges, GNNs can efficiently manage changes in network topology, including the addition or removal of links, without being constrained by the number of nodes. By contrast, node-level tasks may struggle to attain scalability, particularly in scenarios like downlink power allocation. Here, the output dimension at an AP node is typically fixed based on the number of users, limiting flexibility as the network grows. By associating the power allocation variable with the edge, as demonstrated in [46, 47], edge-level GNNs can more effectively adapt to varying network sizes, ensuring scalability.
@@
    Model architecture: We consider the uplink power allocation for a cell-free massive MIMO system, where $M$ APs jointly serve $K$ users simultaneously. The goal is to allocate uplink transmission power for maximizing the minimum user rate across the system. The uplink data rate for each user is adopted from [154]. The global solution may indeed be obtained, but it is computationally expensive, especially for large-scale networks experiencing dynamic user and channel conditions. This makes real-time optimization challenging. Using GNN provides an efficient design alternative.
    As suggested by Peng et al. [45], the cell-free massive MIMO system can be represented as a heterogeneous graph. Therein, the associated max-min fairness problem can be formulated either as an edge-level task or a node-level task and then solved using the edge and node convolution-based methods, respectively. To evaluate the effectiveness of different GNN architectures for this task, we harnessed both models from [45] plus proposed some new ones. Below is a summary of the models considered:
    • Node-GNN: Updates node features, representing users and APs, to allocate power based on node-level characteristics.
    • Hybrid Quantum GNN (Our proposal): The HQGNN uses the GNN to preprocess and generate node embeddings, which are then handled by a deep quantum neural network (DQN) using quantum circuits for power allocation.
    • Edge-GNN: Updates edge features, representing the connections between users and APs, to allocate power based on edge-level characteristics.
    • Graph-GNN: Utilizes global graph pooling to aggregate features across all nodes’ embeddings and then uses the graph embedding vector to predict the power allocation.
@@
    Simulations and results: In this simulation, we consider a cell-free massive MIMO system, where $6$ users are simultaneously served by $30$ access points. The APs and users are uniformly distributed in a square area of size $D\times D$ km2. The path loss model follows a three-slope model based on [155], which accounts for different propagation conditions depending on the AP-UE distance. Only large-scale fading coefficients are used as inputs of the GNNs. We model the system by a bipartite graph associated with two types of nodes, including cellular users and access points. Fig. 8 presents the performance of various GNN architectures, illustrated using the cumulative distribution function (CDF) derived from $10,000$ realizations of a cell-free massive MIMO system.
    The results show that an edge-GNN and node-GNN constitute a pair of models capable of approximating the optimal solution. Briefly, these two models achieve a minimum rate above $1$bps/Hz in about $65$% of the instances. The hybrid quantum GNN, which integrates quantum circuits into the Node-GNN architecture, also performs competitively. Quantitatively, this model tends to achieve $1$bps/Hz in about $64$% of the instances. On average, the Edge-GNN, Node-GNN, and HQGNN models provide data throughputs of $1.0783$, $1.0733$, and $1.0640$ bps/Hz, respectively. By contrast, the optimal solution achieves an average throughput of $1.129$ bps/Hz, slightly outperforming all GNN models, but at the cost of significantly higher computational complexity, particularly as the network size escalates. The graph-GNN performs worst of all GNN models, with less than $40$% of cases reaching a minimum rate of $1.0$ bps/Hz. This poor performance is attributed to its global aggregation of information, which fails to capture the crucial local node and edge interactions necessary for optimizing power allocation in the network. Overall, the results underscore the superiority of node-level GNN models, particularly of those utilizing attention mechanisms for solving the power control problem of cell-free massive MIMO systems. Furthermore, the hybrid quantum GNN offers an intriguing avenue for integrating quantum computing into GNN architectures.
@@ II-3 Challenges and future directions
Again, a significant challenge in modeling GNNs for wireless networks is ensuring scalability, particularly when dealing with networks that have inherent constraints, such as limited node resources or dynamic topology changes. As the graph size grows, the computational cost and memory usage increase, making it challenging to maintain the target performance in real-time. To address this challenge, a promising future direction is to partition the large graph into multiple subgraphs, allowing GNNs to learn locally within each subgraph. The information learned from each subgraph may then be aggregated, ensuring that the model is able to process graphs of varying sizes without losing generalization capabilities. This approach can mitigate scalability issues and maintain performance across different network scales. Moreover, integrating tools from coding theory, such as DE and EXIT charts, offers new opportunities to analyze and improve GNNs in NG-IoT networks. While DE provides convergence guarantees in LDPC decoding by tracking message distributions, its application to GNNs is still developing, particularly in the context of dynamic and heterogeneous graph structures. Similarly, using EXIT charts to monitor mutual information across GNN layers could help identify training bottlenecks like over-smoothing, though more formal metrics and empirical studies are needed. Future research could focus on adapting these tools to analyze GNN convergence, guide model design, and support scalable learning across distributed and partitioned network environments, thereby enabling more interpretable and robust GNN deployment.
Open Question 3: How Can GNNs Be Efficiently Implemented and Optimized for Real-World NG-IoT Systems?
II-1 Background:
NG-IoT systems require real-time decision-making, scalability, and energy efficiency. To efficiently deploy GNNs in NG-IoT systems while meeting these requirements, several key aspects must be considered:
Model Optimization for Edge and Cloud Deployment: Deploying GNN in NG-IoT systems requires careful consideration of computational optimization, energy usage, scalability, and resource-constrainted IoT devices, edge nodes, and cloud platforms:
  • Embedded GNNs for lightweight devices: Embedded GNNs are designed to operate on specialized hardware platforms with limited computational resources, such as AI chips, Field-Programmable Gate Arrays (FPGAs), and Edge Tensor Processing Units (TPUs). Unlike traditional DL models that rely on large-scale GPU or cloud-based processing, embedded GNNs enable on-device learning and inference, reducing the reliance on cloud services and minimizing latency.
  • Model compression techniques for efficient deployment: To address computational constraints, several model compression techniques have been introduced to optimize GNN performance without sacrificing accuracy: 1) Quantization reduces numerical precision in model parameters, storing weights in lower-bit representations such as $8$-bit integers instead of $32$-bit floating points. This technique significantly reduces memory footprint and speeds up inference, making GNNs feasible for real-time applications in NG-IoT systems [53]; Pruning removes unnecessary connections and redundant parameters while preserving the graph structure. By selectively eliminating edges or nodes with minimal contribution to the model’s performance, pruning reduces computational overhead and enhances efficiency. This approach is particularly useful for energy-efficient deployment on edge devices, as demonstrated in recent work on sparse GNNs [54]; Knowledge distillation enables a smaller, compact GNN (student model) to learn from a larger, complex GNN (teacher model). This process allows for transferring knowledge from high-capacity models to lightweight versions while retaining predictive performance.
  • Hardware acceleration for GNN inference: Beyond software-level optimizations, specialized AI hardware has been developed to enhance the execution speed and energy efficiency of GNNs. Hardware accelerators such as Graphics Processing Units (GPUs), Tensor Processing Units (TPUs), Application-Specific Integrated Circuits (ASICs), and neuromorphic chips have been explored for GNN inference, offering improvements in parallelism and computation efficiency. GPUs and TPUs, in particular, have demonstrated substantial speedups in GNN message-passing operations, making them ideal for large-scale graph-based learning in NG-IoT networks [156].
Figure 9: Multi-Modal GNN Framework for NG-IoT Applications.
Multi-modal data processing in NG-IoT networks: Data processing plays a vital role in deploying GNNs in NG-IoT networks, which often involve multimodal data sources, such as sensor readings, network logs, and user interactions, images, or text, as illustrated in Fig. 9. The integration of diverse data modalities presents challenges related to data synchronization and representation [157]. Traditional unimodal processing methods often fail to fully utilize the correlation and complementary information between different modalities. In contrast, GNNs can naturally model the relationships between heterogeneous data sources, capturing their interdependencies in complex IoT environments [56, 57]. Fig. 9 illustrates how heterogeneous data sources, such as images, temperature, humidity, GPS, voice, and text, are processed through dedicated multi-modal tools (e.g., image processing, text analysis, and sensor embeddings) to generate graph-structured data. A multi-modal GNN then learns from these fused graph representations to support various NG-IoT tasks, including energy-efficient routing, resource allocation, anomaly detection, smart monitoring, device failure prediction, and fault detection.
However, ensuring the scalability and efficiency of multimodal GNNs in real-world NG-IoT systems remains an open challenge, requiring further research into data pre-processing, representation learning, and computational optimization techniques.
Security, Privacy, and Scalability Challenges: As NG-IoT applications involve large-scale, interconnected devices handling sensitive data, ensuring the privacy of IoT users and the security of IoT applications is critical. Indeed, IoT networks are inherently vulnerable to adversarial attacks and inference manipulation, hence necessitating robust security mechanisms. To address these concerns, encryption-based training, differential privacy techniques, and secure multi-party computation can enhance security for GNN deployment. Given the importance of these challenges, we provide an in-depth discussion on adversarial attacks and defense mechanisms for GNN-based NG-IoT networks in Section IV, analyzing different attack strategies and mitigation techniques.
Socio-Economic Dimensions of NG-IoT Applications: The deployment of GNNs in NG-IoT networks introduces several socio-economic considerations, including cost, sustainability, and accessibility.
  • Cost consideration: Deploying GNNs in NG-IoT networks involves significant costs, including training, deployment, and scalability. The training cost includes computational resources, energy, and hardware costs [58]. Deploying GNNs involves infrastructure costs, such as purchasing and maintaining edge devices (e.g., NVIDIA Jetson, Raspberry Pi) or cloud-based resources (e.g., AWS, Google Cloud). Operational costs, such as energy consumption and maintenance, also add to the overall expense [59]. Meanwhile, scaling GNN-based solutions to large IoT networks requires additional investments in hardware, software, and network infrastructure [60].
  • Sustainability: Training and deploying GNNs on IoT devices, especially edge devices, can be energy-intensive. For example, training even lightweight GNN models on edge devices can strain battery life and increase operational costs [61]. Besides, the energy consumption of GNNs contributes to carbon emissions, particularly when non-renewable energy sources are used [61]. Accordingly, this poses sustainable challenges limiting the adoption of GNNs in NG-IoT networks, especially in regions with strict environmental regulations or limited access to renewable energy. Green AI initiatives, such as energy-efficient algorithms and renewable energy-powered data centers, are critical to addressing these challenges
II-2 State-of-the-arts
Model Optimization for Edge and Cloud Deployment: To streamline the deployment of GNNs in NG-IoT networks, various studies have explored different key aspects. In attempt to embed GNN, Zhang et al. [52] have explored how (FPGAs) can accelerate message-passing operations in GNNs, leading to faster inference times while maintaining model accuracy. Besides, the authors also summarizes the challenges that FPGA-based GNNs accelerators need to address. Meanwhile, Wang et al. [53] proposes a quantization-based approach to address scalability challenges in GNNs. By compressing models with low-bit quantization and controlling message propagation, it reduces computational and memory costs while maintaining accuracy. The method achieves significant speedups (5.11×5.$1$ for INT2, 4.70×4.70$\times$4.70 × for INT4), offering a cost-effective solution for deploying GNNs in NG-IoT networks. Xie et al. [54] proposed Graph sparsification and network pruning (GASSIP), a novel framework for lightweight graph neural architecture search (GNAS) tailored for resource-constrained scenarios. By combining operation-pruned architecture search and curriculum graph sparsification, GASSIP efficiently identifies optimal lightweight GNNs with reduced model parameters and sparser graphs. Extensive experiments show that GASSIP achieves comparable or superior performance with significantly fewer resources, making it highly relevant for cost-effective GNN deployment in NG-IoT applications. In [55], Wu et al. introduced Prototype-Guided Knowledge Distillation (PGKD), a novel method for distilling high-accuracy GNNs into low-latency MLPs without relying on graph edges. By leveraging class prototypes and structure-aware distillation losses, PGKD effectively transfers graph structural knowledge from GNNs to MLPs, enabling edge-free learning.
Multi-modal data processing in NG-IoT networks: Li et al. [56] proposed GNNMR, a multi-modal recommendation framework that integrates GNNs with deep mutual learning to address modality bias. By training separate GNNs on uni-modal user-item graphs and using mutual knowledge distillation, GNNMR synchronizes latent semantic relationships across modalities, improving multi-modal embeddings. It outperforms existing methods in Top-K recommendation tasks, making it suitable for NG-IoT applications requiring efficient user-item interaction modeling. Similarly, Zhang et al. [57] introduced a GNN framework for multimodal data integration, leveraging feature, decision, and deep fusion techniques to achieve high accuracy ($98$%), recall ($86.9$%), and F1 score ($0.964$). This approach demonstrates robust performance in complex scenarios, highlighting the potential of multimodal GNNs for NG-IoT applications.
Cost consideration: Recent studies have explored various strategies to address the socio-economic challenges of deploying GNNs in NG-IoT systems, particularly focusing on reducing deployment and operational costs, enhancing scalability, and improving energy efficiency. Shuvo et al. [58] emphasized the cost-efficiency benefits of deploying deep learning models on edge devices in NG-IoT systems. While cloud-based HPC clusters provide high computational capacity, they incur high transmission costs, latency, and privacy concerns. In contrast, edge computing with devices like IoT sensors and wearables can offer real-time processing at reduced operational costs. Duc et al. [59] addressed operational costs in edge-cloud deployments by proposing ML-based resource provisioning frameworks. Their study examined workload characterization, elastic resource management, and component placement to minimize energy consumption and infrastructure usage, which are key challenges in large-scale NG-IoT applications. In [60], Murshed et al. provided a comprehensive overview of deploying ML models at the edge to reduce latency and offloading costs. They reviewed compression techniques, frameworks, and hardware accelerators suited for real-time ML applications in constrained environments, supporting scalable and low-latency NG-IoT deployment. In parallel, Strubell et al. [61] quantified the environmental costs of training large AI models and advocated for sustainable AI practices. Their findings support the shift toward low-power, energy-efficient GNN deployment strategies in NG-IoT, aligning with green AI initiatives to reduce carbon footprint and promote equitable access to AI technologies.
II-3 Challenges and future direction
Despite recent advances, several challenges remain in deploying GNNs efficiently within real-world NG-IoT systems. One of the primary issues is the computational and energy constraints of edge devices, such as IoT sensors and mobile nodes. These resource limitations demand advanced model optimization techniques, including pruning, quantization, and knowledge distillation, alongside hardware-software co-design strategies that tailor GNN architectures to run efficiently on FPGAs, TPUs, and neuromorphic chips. Furthermore, as NG-IoT systems scale in size and complexity, ensuring real-time inference and learning becomes increasingly difficult. This necessitates scalable training frameworks, decentralized and federated GNN models, and intelligent graph partitioning and sampling techniques.
Another significant challenge is the integration and processing of multimodal data (e.g., sensor readings, logs, and user interactions), which requires effective fusion strategies and synchronized graph representations. In addition, from the socio-economic perspective, cost-efficiency and sustainability remain key concerns, which hinder wide-scale adoption. Hence, promoting energy-efficient algorithms, carbon-aware deployment strategies, and global standardization efforts will be essential for the practical, scalable, and equitable integration of GNNs in NG-IoT applications.